OpenStudy (anonymous):

A large inlet pipe can fill a tank 16 minutes faster than a smaller pipe. When both inlets are used, the tank fills in 6 minutes. How long does it take the smaller inlet pipe working alone to fill the tank?

OpenStudy (lgbasallote):

let x be the time it takes for the smaller pipe to fill the tank (alone) therefore it takes x+ 16 for the bigger pipe now we use the formula \[\frac{t}{a} +\frac{t}{b} = 1\] where a and b are the individual time and t is the total time it says the total time here is 6 minutes so \[\frac{6}{x} + \frac{6}{x+ 16} = 1\] now solve for x

OpenStudy (anonymous):

let rate of work for small pipe be 1/p then rate of work for big pipe is 1/(p+16) then if they work for 6 min the work is done that is ,  p cannot be negative so p = 8 min